OpenStudy (anonymous):

1. Solve by completing the square 4r^2 - 28r = -9 2. Use the discriminant to find the number of unique real solutions. 3. Solve with the quadratic formula.

OpenStudy (anonymous):

wat did u get for number 1?

OpenStudy (anonymous):

I didn't solve it, not sure how.

OpenStudy (anonymous):

well first we rearrange the formula to get 4r^2 - 28r = -9 4r^2 - 28r + 9 = 0 and from here u factorize the equation

OpenStudy (anonymous):

do u know how to factorize an equation or do u want me to teach u how?

OpenStudy (anonymous):

not sure how, please teach me. I'm horrible with math

OpenStudy (anonymous):

ok ill do it and ill explan wat i did after i finish

OpenStudy (anonymous):

Alright, thank you. I appreciate it.

OpenStudy (anonymous):

4r^2 - 28r + 9 = 0 4(r^2 - 7r) +9 \[4(r - \frac{ 7 }{ 2 })^{2} + 9 - 4(\frac{ 7 }{ 2 })^{2}\] \[4(r - \frac{ 7 }{ 2 })^{2} + 9 - 4(\frac{ 49 }{ 4 })\] \[4(r - \frac{ 7 }{ 2 })^{2} + 9 - 49 \] \[4(r - \frac{ 7 }{ 2 })^{2} - 40\]

OpenStudy (anonymous):

ok i need u to be here and focus, u there?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

sry im back my internet was giving trouble, u there again?